OpenStudy (smileyxl3):

If your starting salary were $35,000 and you received a 6% increase at the end of every year for 15 years, what would be the total amount, in dollars, you would have earned over the first 16 years that you worked?

OpenStudy (smileyxl3):

Is this the equation? 35000 * (1.06)^15

OpenStudy (anonymous):

I would do 35000* 6% and then multiply the answer by fifteen and add that answer to 35000.

OpenStudy (anonymous):

I'll do both yours and mine and see if the answer is the same.

OpenStudy (smileyxl3):

lol your way is 66500...

OpenStudy (anonymous):

hmmm.... two very different answers. my guess is you were right. :P

OpenStudy (anonymous):

Sometimes I'm good at math and sometimes... welllll....

OpenStudy (smileyxl3):

Ok kel, thanks for your help

OpenStudy (smileyxl3):

@FabiodaSilva What do you think?

OpenStudy (anonymous):

I did it like 35000 x 0.06 = 2100 2100 x 16 = 33600 35000 x 16 = 560000 560000 + 33600 = 593600 That's how much you'd earn at the end of 16 years

OpenStudy (anonymous):

I think anyways

OpenStudy (smileyxl3):

Oops I forgot to add the last year to my original equation... should be 83879.54 + 35000... I think

OpenStudy (anonymous):

It's long but bare with me. 35000 x 0.06 = 2100 35000 + 2100 = 37100 (First year) 37100 x 0.06 = 2226 37100 + 2226 = 39326 (Second year) Etc etc. There is probably a faster way to do this, but this is the only way i can think of.

OpenStudy (smileyxl3):

Ok, turns out the right answer was 814659. How is that the answer though? @Compassionate

OpenStudy (anonymous):

I know this is old but I just found out the equation so for future reference for when people find this thread later on use the formula Sn=A1(1-R^n)/1-R